# Cluster algebras and quantum cohomology rings: A-type

### v1

@@ 5.1.1 Abelian/nonabelian correspondence for cohomology rings
We first review the abelian/nonabelian correspondence for cohomology rings, and refer to [ESm89, Mar00].
@@
Let $\Phi=\{\rho\}$ be the set of positive roots of $G$, and $\{\mathbb{C}_{\rho}\}_{\rho\in\Phi}$ the one dimensional representations of $T$ with weights $\rho$. These one dimensional representations define line bundles on $V\sslash_{\theta}T$, which are denoted by $L_{\rho}=V\times_{T}\mathbb{C}_{\rho}$. Denote their equivariant first Chern classes by $c^{S}_{1}(L_{\rho})\in H_{S}^{*}(V\sslash_{\theta}T)$. Define $\omega:=\prod_{\rho\in\Phi}c^{S}_{1}(L_{\rho})$. Then we have the following facts [CFKS08, Section 3.1]:
@@
    $0\rightarrow\ker(\omega)\rightarrow H^{*}_{S}(V\sslash_{\theta}T)^{W}\xrightarrow[]{(p^{*})^{-1}\circ a^{*}}H^{*}_{S}(V\sslash_{\theta}G)\rightarrow 0,$ (5.8)
@@ 5.1.2 Abelian/nonabelian correspondence for quantum cohomology rings
There is a quantum version of (5.8) which we review by following the work [CFKS08]. Let $U$ be a subspace in $\subset H^{*}(V\sslash_{\theta}T)$ which is a lift of $H^{*}(V\sslash_{\theta}G)$. Denote the lifting of an element $\gamma$ by $\tilde{\gamma}$. Let $\mathbf{t}=\{t_{i}\}$ be the coordinates on $H^{*}(X\sslash_{\theta}G)$ by choosing a basis, and $\tilde{\mathbf{t}}=\{\tilde{t}_{i}\}$ be the coordinates on $U$ with respect to the lifted basis. Let $(\tilde{\mathbf{t}},\mathbf{y})$ be an extension of $\tilde{\mathbf{t}}$ to the entire space $H^{*}(X\sslash_{\theta}T)$.
@@
which takes into consideration that there are more curve classes in $V\sslash_{\theta}T$. We denote by “$*_{g}$” the big quantum product on $V\sslash_{\theta}T$ with the Novikov variables being specialized via the map $\pi$ or on $V\sslash_{\theta}G$ depending on where the cohomology classes are. The abelian/nonabelian correspondence for quantum cohomology ring is as follows [CFKS08, Conjecture 3.7.1].
@@
([CFKS08, Conjecture 3.7.1]) Given $\gamma,\gamma^{\prime}\in H^{*}(V\sslash_{\theta}G)$, there are classes $\zeta,\zeta^{\prime}\in U\otimes_{\mathbb{C}}N(V\sslash_{\theta}G)[[\tilde{\mathbf{t}}]]$, uniquely determined by relations $\zeta*_{g}\omega=\tilde{\gamma}\cup\omega$ and $\zeta^{\prime}*_{g}\omega=\tilde{\gamma}^{\prime}\cup\omega$. There is an equality
$(\widetilde{(\gamma*_{g}\gamma^{\prime})}\cup\omega)(\mathbf{t})=\zeta*_{g}\zeta^{\prime}*_{g}\omega(\tilde{\mathbf{t}},0)=(\zeta*_{g}(\widetilde{\zeta^{\prime}}\cup\omega))(\tilde{\mathbf{t}},0),$ (5.9)
@@
Furthermore, (5.9) is proved in the case of flag varieties $Fl$ in [CFKS08, Theorem 4.1.1], and is proved under some assumptions in [Web23, Corrolary 6.4.5]. See [GK24] for an interesting application of abelian/nonabelian correspondence. We will only use the restriction of (5.9) to small quantum cohomology ring in our case $V\sslash_{\theta}G=Fl=Fl(N_{1},\cdots,N_{n+1})$. We will use * to denote the small quantum product in $V\sslash_{\theta}G$ or $V\sslash_{\theta}T$ with Novikov variables specialized via the map $\pi$.
@@
$(\widetilde{\gamma*\gamma^{\prime}})\cup\omega=\xi*\xi^{\prime}*\omega=\xi*(\widetilde{\gamma^{\prime}}\cup\omega),$ (5.10)
Notice that $\gamma*\gamma^{\prime}$ is uniquely determined in (5.10) according to the exact sequence (5.7).
@@
The variables change $\widetilde{t}=\widetilde{t}(t)$ in (5.9) is the only thing we need to care about. This change is trivial after restricted to the small phase subspace since the abelian/nonabelian correspondence for small J-functions holds for $Fl$ without parameter change , see [CFKS08, Theorem 4.36] or [GK24, Section 4.2]. ∎
5.2 Quantum cohomology of the abelianization of $Fl(N_{1},\cdots,N_{n+1})$
Consider the flag variety $Fl:=Fl(N_{1},\ldots,N_{n+1})$, which can be written as $V\sslash_{\theta}G$ as in Example 2.14. The gauge group is $G=\prod_{i=1}^{n}GL(N_{i})$, the maximal torus is $T=\prod_{i=1}^{n}(\mathbb{C}^{*})^{N_{i}}$, and the weyl group $W=\prod_{i=1}^{n}S_{N_{i}}$ where $S_{N_{i}}$ is the permutation group of $N_{i}$ elements. The abelianization $V\sslash_{\theta}T$ is a toric variety, which can be viewed as a tower of product of projective spaces, $(\mathbb{P}^{N_{2}-1})^{N_{1}}\rightarrow(\mathbb{P}^{N_{3}-1})^{N_{2}}\rightarrow\ldots\rightarrow(\mathbb{P}^{N_{n+1}-1})^{N_{n}}$. Let $[\ell_{k,i}]$ denote the rational curve class lying inside the $i$-th $\mathbb{P}^{N_{k+1}-1}$-fiber of the $k$-th tower $(\mathbb{P}^{N_{k+1}-1})^{N_{k}}$. Then $N(V\sslash_{\theta}T)=\bigoplus_{k,i}\mathbb{Z}_{\geq 0}[\ell_{k,i}]$. Set the Novikov variable $q_{k,i}:=q^{[\ell_{k,i}]}$.
@@ 5.2 Quantum cohomology of the abelianization of $Fl(N_{1},\cdots,N_{n+1})$
Consider characters $\chi_{j}^{(i)}:T\rightarrow\mathbb{C}^{*}$ which maps $(\vec{t}^{i})_{i=1}^{n}\in T$ to the $j$-th component of the $\vec{t}^{i}=\{t^{i}_{1},\cdots,t^{i}_{N_{i}}\}\in(\mathbb{C}^{*})^{N_{i}}$. These characters define a set of line bundles $L_{\chi^{(i)}_{j}}$ over $V\sslash_{\theta}T$. Denote
$x^{(i)}_{j}=c_{1}^{S}(L_{\chi^{(i)}_{j}})\in H^{*}_{S}(V\sslash_{\theta}T),\text{ and }\vec{x}^{(i)}=\{x^{(i)}_{1},\cdots x^{(i)}_{N_{i}}\}$ (5.11)
Actually, for each $i=1,\cdots,n,j=1,\cdots,N_{i}$, the non-equivariant limit of $x^{(i)}_{j}$ is the hyperplane class of the $j$-th projective space of $(\mathbb{P}^{N_{i+1}})^{N_{i}}$.
@@
$\prod_{l=1}^{N_{i+1}}(x^{(i)}_{j}-x^{(i+1)}_{l})=0,\,\,j=1,\ldots,N_{i};\,\,i=1,\ldots,n,$ (5.12)
@@
$H^{*}_{S}(V\sslash_{\theta}T)=\mathbb{Q}[\{x^{(i)}_{j}\}_{i=1,\cdots,n+1,j=1,\cdots,N_{i}}]/I^{ab}.$ (5.13)
The quantum cohomology rings of toric varieties have been investigated by Batyrev [Bat93]. We can directly write down the quantum cohomology of $V\sslash_{\theta}T$ as follows.
@@
$\left\{\prod_{k=1}^{n}\prod_{j=1}^{N_{k}}\left(x^{(k)}_{j}\right)^{a_{kj}}\big{|}0\leq a_{kj}\leq N_{k+1}-1\right\}$ (5.14)
@@
$=-\sum_{l=0}^{N_{k+1}-1}e_{N_{k+1}-l}(-\vec{x}^{(k+1)})\left(x^{(k)}_{i}\right)^{l}+q_{k,i}\prod_{l=1}^{N_{k-1}}\left(x^{(k-1)}_{l}-x^{(k)}_{i}\right)$ (5.15)
@@
From (5.12) and (5.6), one can find that $*|_{q\rightarrow 0}=\cup$. For any $\alpha,\beta\in H^{*}_{S}(V\sslash_{\theta}T)$, we call $\alpha\cup\beta$ the classical part and $\alpha*\beta-\alpha\cup\beta$ the quantum correction of $\alpha*\beta$. Then we know from (5.6) that the quantum correction of $\left(x^{(k)}_{i}\right)^{N_{k+1}-1}*x^{(k)}_{i}$ is $q_{k,i}\prod_{l=1}^{N_{k-1}}\left(x^{(k-1)}_{l}-x^{(k)}_{i}\right)$, and the classical part is $-\sum_{l=0}^{N_{k+1}-1}e_{N_{k+1}-l}(-\vec{x}^{(k+1)})\left(x^{(k)}_{i}\right)^{l}$.
From Proposition 5.6, we can easily get that the quantum correction of $\alpha*\beta$ is nontrivial only if the sum of the degree of $x^{(k)}_{i}$ in $\alpha$ and $\beta$ $\geq N_{k+1}$ for some $k$ and $i$.
5.3 Quantum cohomology ring of $Fl(N_{1},\ldots,N_{n+1})$
@@
$\prod_{1\leq i<j\leq N}(x_{i}-x_{j})=\det\begin{bmatrix}x_{1}^{N-1}&x_{2}^{N-1}&\ldots&x_{N}^{N-1}\\ x_{1}^{N-2}&x_{2}^{N-2}&\ldots&x_{N}^{N-2}\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\ldots&1\end{bmatrix}.$ (5.16)
@@
$s_{\eta}(x_{1},\ldots,x_{N})\prod_{1\leq i<j\leq N}(x_{i}-x_{j})=\det\begin{bmatrix}x_{1}^{\eta_{1}+N-1}&x_{2}^{\eta_{1}+N-1}&\cdots&x_{N}^{\eta_{1}+N-1}\\ x_{1}^{\eta_{2}+N-2}&x_{2}^{\eta_{2}+N-2}&\cdots&x_{N}^{\eta_{2}+N-2}\\ \vdots&\vdots&\ddots&\vdots\\ x_{1}^{\eta_{N}}&x_{2}^{\eta_{N}}&\cdots&x_{N}^{\eta_{N}}\end{bmatrix}.$ (5.17)
@@
$h_{i}(x_{1},\cdots,x_{N})=\sum_{j_{1}+\ldots+j_{N}=i}\prod_{l=1}^{N}{x_{l}^{j_{l}}}.$ (5.18)
We denote $\omega_{m}=\prod_{1\leq i<j\leq N}(x_{i}^{(m)}-x_{j}^{(m)})$ where $x^{(m)}_{i}$ are Chern roots of line bundles defined in (5.11).
Consider the equivariant cohomology ring of $Fl$. By the short exact sequence (5.1), the equivariant cohomology ring of $Fl$ is
$H^{*}_{S}(Fl)\cong\mathbb{Q}[\{x^{(i)}_{j}\}_{i=1,\cdots,n+1,j=1,\cdots,N_{i}}]^{W}/I$ (5.19)
where the ideal $I$ is generated by elements $\{f\in\mathbb{Q}[\lambda_{1},\ldots,\lambda_{N_{n+1}}][\{x^{(i)}_{j}\}_{i=1,\cdots,n+1,j=1,\cdots,N_{i}}]^{W}\big{|}\,f\cup\omega\in I^{ab}\}$ where $\omega$ can be written as
$\omega=\prod_{k=1}^{n}\omega_{k}=\prod_{k=1}^{n}\prod_{1\leq i<j\leq N_{k}}(x^{(k)}_{i}-x^{(k)}_{j}).$ (5.20)
@@
$\widetilde{c_{t}^{S}(S_{k})}=\prod_{j=1}^{N_{k}}(t-x^{(k)}_{j})=\sum_{l=0}^{N_{k}}t^{N_{k}-l}e_{l}(-\vec{x}^{(k)}).$ (5.21)
For $k<l$, let $S_{l}/S_{k}$ be the quotient bundle defined by the following exact sequence
$0\rightarrow S_{k}\rightarrow S_{l}\rightarrow S_{l}/S_{k}\rightarrow 0$ (5.22)
Then in $H^{*}_{S}(Fl)$, the Chern polynomial of $S_{l}/S_{k}$ satisfies the following relation
$c_{t}^{S}(S_{k})\cup c_{t}^{S}(S_{l}/S_{k})=c_{t}^{S}(S_{l}).$ (5.23)
@@
$\prod_{j=1}^{N_{l}}\left(x^{(k)}_{i}-x^{(l)}_{j}\right)\prod_{m=k+1}^{l-1}\omega_{m}=0,\,\,i=1,2,\ldots,N_{k}.$ (5.24)
@@
We use an induction argument on the difference $l-k$. For $l-k=1$, the equation (5.24) holds by (5.12). We assume (5.24) holds for $l-k\leq p-1$. For $l-k=p$, we consider
$A:=\det\begin{bmatrix}\prod_{j=1}^{N_{l}}\left(x^{(k)}_{i}-x^{(l)}_{j}\right)&\prod_{j=1}^{N_{l}}\left(x_{1}^{(l-1)}-x^{(l)}_{j}\right)&\cdots&\prod_{j=1}^{N_{l}}\left(x_{N_{l-1}}^{(l-1)}-x^{(l)}_{j}\right)\\ \left(x_{i}^{(k)}\right)^{N_{l-1}-1}&\left(x_{1}^{(l-1)}\right)^{N_{l-1}-1}&\cdots&\left(x_{N_{l-1}}^{(l-1)}\right)^{N_{l-1}-1}\\ \vdots&\vdots&\ddots&\vdots\\ x_{i}^{(k)}&x_{1}^{(l-1)}&\cdots&x_{N_{l-1}}^{(l-1)}\\ 1&1&\cdots&1\\ \end{bmatrix}\cdot\prod_{m=k+1}^{l-2}\omega_{m}$ (5.25)
@@
$A=\sum_{a=1}^{N_{l}}e_{a}(-\vec{x}^{(l)})h_{N_{l}-N_{l-1}-a}(x^{(k)}_{i},\vec{x}^{(l-1)})\prod_{j=1}^{N_{l-1}}\left(x^{(k)}_{i}-x^{(l-1)}_{j}\right)\prod_{m=k+1}^{l-1}\omega_{m}$ (5.26)
@@
On the other hand, expanding (5.25) along the first row, by (5.12), (5.16) we have
$A=\prod_{j=1}^{N_{l}}\left(x^{(k)}_{i}-x^{(l)}_{j}\right)\prod_{m=k+1}^{l-1}\omega_{m}$ (5.27)
So (5.24) holds for $l-k=p$. ∎
@@
$\widetilde{c_{t}^{S}(S_{l}/S_{k})}=\sum_{a=0}^{N_{l}-N_{k}}t^{a}\left[\sum_{b=0}^{N_{l}-N_{k}-a}e_{b}(-\vec{x}^{(l)})h_{N_{l}-N_{k}-a-b}(\vec{x}^{(k)})\right].$ (5.28)
@@
By (5.8) and (5.23), it suffices to check
$\widetilde{c_{t}^{S}(S_{k})}\cup\sum_{a=0}^{N_{l}-N_{k}}t^{a}\left[\sum_{b=0}^{N_{l}-N_{k}-a}e_{b}(-\vec{x}^{(l)})h_{N_{l}-N_{k}-a-b}(\vec{x}^{(k)})\right]\cup\omega=\widetilde{c_{t}^{S}(S_{l})}\cup\omega.$ (5.29)

### v2

@@ 5.1.1 Abelian/nonabelian correspondence for cohomology rings
We first review the abelian/nonabelian correspondence for cohomology rings, and refer to [Ab_nab:EG, Ab_nab:Mar].
@@
Let $\Phi=\{\rho\}$ be the set of positive roots of $G$, and $\{\mathbb{C}_{\rho}\}_{\rho\in\Phi}$ the one dimensional representations of $T$ with weights $\rho$. These one dimensional representations define line bundles on $V\sslash_{\theta}T$, which are denoted by $L_{\rho}=V\times_{T}\mathbb{C}_{\rho}$. Denote their equivariant first Chern classes by $c^{S}_{1}(L_{\rho})\in H_{S}^{*}(V\sslash_{\theta}T)$. Define $\omega:=\prod_{\rho\in\Phi}c^{S}_{1}(L_{\rho})$. Then we have the following facts [abelian/nonabelian:CKB, Section 3.1]:
@@
    (37) $0\rightarrow\ker(\omega)\rightarrow H^{*}_{S}(V\sslash_{\theta}T)^{W}\xrightarrow[]{(p^{*})^{-1}\circ a^{*}}H^{*}_{S}(V\sslash_{\theta}G)\rightarrow 0,$
@@ 5.1.2 Abelian/nonabelian correspondence for quantum cohomology rings
There is a quantum version of (37) which we review by following the work [abelian/nonabelian:CKB]. Let $U$ be a subspace in $H^{*}(V\sslash_{\theta}T)$ which is a lift of $H^{*}(V\sslash_{\theta}G)$. Denote the lifting of an element $\gamma$ by $\tilde{\gamma}$. Let $\mathbf{t}=\{t_{i}\}$ be the coordinates on $H^{*}(X\sslash_{\theta}G)$ by choosing a basis, and $\tilde{\mathbf{t}}=\{\tilde{t}_{i}\}$ be the coordinates on $U$ with respect to the lifted basis. Let $(\tilde{\mathbf{t}},\mathbf{y})$ be an extension of $\tilde{\mathbf{t}}$ to the entire space $H^{*}(X\sslash_{\theta}T)$.
@@
which takes into consideration that there are more curve classes in $V\sslash_{\theta}T$. We denote by “$*_{g}$” the big quantum product on $V\sslash_{\theta}T$ with the Novikov variables being specialized via the map $\pi$ or on $V\sslash_{\theta}G$ depending on where the cohomology classes are. The abelian/nonabelian correspondence for quantum cohomology ring is as follows [abelian/nonabelian:CKB, Conjecture 3.7.1].
@@
([abelian/nonabelian:CKB, Conjecture 3.7.1]) Given $\gamma,\gamma^{\prime}\in H^{*}(V\sslash_{\theta}G)$, there are classes $\zeta,\zeta^{\prime}\in U\otimes_{\mathbb{C}}N(V\sslash_{\theta}G)[[\tilde{\mathbf{t}}]]$, uniquely determined by relations $\zeta*_{g}\omega=\tilde{\gamma}\cup\omega$ and $\zeta^{\prime}*_{g}\omega=\tilde{\gamma}^{\prime}\cup\omega$. There is an equality
(38) $(\widetilde{(\gamma*_{g}\gamma^{\prime})}\cup\omega)(\mathbf{t})=\zeta*_{g}\zeta^{\prime}*_{g}\omega(\tilde{\mathbf{t}},0)=(\zeta*_{g}(\widetilde{\zeta^{\prime}}\cup\omega))(\tilde{\mathbf{t}},0),$
@@
The conjecture 38 is proved in the case of flag varieties $Fl$ in [abelian/nonabelian:CKB, Theorem 4.1.1], and is proved under the assumptions that the GIT quotient $V\sslash_{\theta}G$ is Fano of index greater than or equal to two and the equivariant cohomology ring $H^{*}_{R}(V\sslash_{\theta}G)$ is generated by divisor classes after localization in [abeliannonabelian:Webb, Corrolary 6.4.5] and [abelian/nonabelian:CKB, Theorem 4.3.6]. See [rimhookrule] for an interesting application of abelian/nonabelian correspondence. We will only use the restriction of (38) to small quantum cohomology ring in our case $V\sslash_{\theta}G=Fl=Fl(N_{1},\cdots,N_{n+1})$. We will use * to denote the small quantum products in $V\sslash_{\theta}G$ and $V\sslash_{\theta}T$ with Novikov variables being specialized via the map $\pi$.
@@
(39) $(\widetilde{\gamma*\gamma^{\prime}})\cup\omega=\xi*\xi^{\prime}*\omega=\xi*(\widetilde{\gamma^{\prime}}\cup\omega),$
Notice that $\gamma*\gamma^{\prime}$ is uniquely determined in (39) according to the exact sequence (36).
@@
The variables change $\widetilde{t}=\widetilde{t}(t)$ in (38) is the only thing we need to care about. This change is trivial after being restricted to the small phase subspace since the abelian/nonabelian correspondence for small J-functions holds for $V\sslash_{\theta}G$ without parameter change, see [abelian/nonabelian:CKB, Theorem 4.3.6] or [rimhookrule, Section 4.2]. ∎
5.2 Quantum cohomology of the abelianization of a flag variety
Consider the flag variety $Fl:=Fl(N_{1},\ldots,N_{n+1})$, which can be written as $V\sslash_{\theta}G$ as in Example 2.14. The gauge group is $G=\prod_{i=1}^{n}GL(N_{i})$, the maximal torus is $T=\prod_{i=1}^{n}(\mathbb{C}^{*})^{N_{i}}$, and the Weyl group $W=\prod_{i=1}^{n}S_{N_{i}}$ where $S_{N_{i}}$ is the permutation group of $N_{i}$ elements. The abelianization $V\sslash_{\theta}T$ is a toric variety, which can be viewed as a tower of products of projective spaces, $(\mathbb{P}^{N_{2}-1})^{N_{1}}\rightarrow(\mathbb{P}^{N_{3}-1})^{N_{2}}\rightarrow\ldots\rightarrow(\mathbb{P}^{N_{n+1}-1})^{N_{n}}$. Let $[\ell_{k,i}]$ denote the rational curve class lying inside the $i$-th $\mathbb{P}^{N_{k+1}-1}$-fiber of the $k$-th tower $(\mathbb{P}^{N_{k+1}-1})^{N_{k}}$. Then $N(V\sslash_{\theta}T)=\bigoplus_{k,i}\mathbb{Z}_{\geq 0}[\ell_{k,i}]$. Set the Novikov variable $q_{k,i}:=q^{[\ell_{k,i}]}$.
@@ 5.2 Quantum cohomology of the abelianization of a flag variety
Consider characters $\chi_{j}^{(i)}:T\rightarrow\mathbb{C}^{*}$ which map $(\vec{t}^{i})_{i=1}^{n}\in T$ to the $j$-th component of the $\vec{t}^{i}=\{t^{i}_{1},\cdots,t^{i}_{N_{i}}\}\in(\mathbb{C}^{*})^{N_{i}}$. These characters define a set of line bundles $L_{\chi^{(i)}_{j}}$ over $V\sslash_{\theta}T$. Denote
(40) $x^{(i)}_{j}=c_{1}^{S}(L_{\chi^{(i)}_{j}})\in H^{*}_{S}(V\sslash_{\theta}T),\text{ and }\vec{x}^{(i)}=\{x^{(i)}_{1},\cdots x^{(i)}_{N_{i}}\}\,.$
Actually, for each $i=1,\cdots,n,j=1,\cdots,N_{i}$, the non-equivariant limit of $x^{(i)}_{j}$ is the hyperplane class of the $j$-th projective space of $(\mathbb{P}^{N_{i+1}-1})^{N_{i}}$.
@@
(41) $\prod_{l=1}^{N_{i+1}}(x^{(i)}_{j}-x^{(i+1)}_{l})=0,\,\,j=1,\ldots,N_{i};\,\,i=1,\ldots,n,$
@@
$H^{*}_{S}(V\sslash_{\theta}T)=\mathbb{Q}[\{x^{(i)}_{j}\}_{i=1,\cdots,n+1,j=1,\cdots,N_{i}}]/I^{ab}.$
The quantum cohomology rings of toric varieties have been investigated by Batyrev [toric]. We can directly write down the quantum cohomology of $V\sslash_{\theta}T$ as follows.
@@
(42) $\left\{\prod_{k=1}^{n}\prod_{j=1}^{N_{k}}\left(x^{(k)}_{j}\right)^{a_{kj}}\big{|}0\leq a_{kj}\leq N_{k+1}-1\right\}$
@@
(43) $=-\sum_{l=0}^{N_{k+1}-1}e_{N_{k+1}-l}(-\vec{x}^{(k+1)})\left(x^{(k)}_{i}\right)^{l}+q_{k,i}\prod_{l=1}^{N_{k-1}}\left(x^{(k-1)}_{l}-x^{(k)}_{i}\right)\,.$
@@
From (41) and (5.6), one can find that $*|_{q\rightarrow 0}=\cup$. For any $\alpha,\beta\in H^{*}_{S}(V\sslash_{\theta}T)$, we call $\alpha\cup\beta$ the classical part and $\alpha*\beta-\alpha\cup\beta$ the quantum correction of $\alpha*\beta$. Then we know from (5.6) that the quantum correction of $\left(x^{(k)}_{i}\right)^{N_{k+1}-1}*x^{(k)}_{i}$ is $q_{k,i}\prod_{l=1}^{N_{k-1}}\left(x^{(k-1)}_{l}-x^{(k)}_{i}\right)$, and the classical part is $-\sum_{l=0}^{N_{k+1}-1}e_{N_{k+1}-l}(-\vec{x}^{(k+1)})\left(x^{(k)}_{i}\right)^{l}$.
From Proposition 5.6, we can easily get that the quantum correction of $\alpha*\beta$ is nontrivial only if the sum of the degree of $x^{(k)}_{i}$ in $\alpha$ and $\beta$ is greater than or equal to $N_{k+1}$ for some $k$ and $i$.
5.3 Quantum cohomology rings of flag varieties
@@
(44) $\prod_{1\leq i<j\leq N}(x_{i}-x_{j})=\det\begin{bmatrix}x_{1}^{N-1}&x_{2}^{N-1}&\ldots&x_{N}^{N-1}\\ x_{1}^{N-2}&x_{2}^{N-2}&\ldots&x_{N}^{N-2}\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\ldots&1\end{bmatrix}.$
@@
(45) $s_{\eta}(x_{1},\ldots,x_{N})\prod_{1\leq i<j\leq N}(x_{i}-x_{j})=\det\begin{bmatrix}x_{1}^{\eta_{1}+N-1}&x_{2}^{\eta_{1}+N-1}&\cdots&x_{N}^{\eta_{1}+N-1}\\ x_{1}^{\eta_{2}+N-2}&x_{2}^{\eta_{2}+N-2}&\cdots&x_{N}^{\eta_{2}+N-2}\\ \vdots&\vdots&\ddots&\vdots\\ x_{1}^{\eta_{N}}&x_{2}^{\eta_{N}}&\cdots&x_{N}^{\eta_{N}}\end{bmatrix}.$
@@
$h_{i}(x_{1},\cdots,x_{N})=\sum_{j_{1}+\ldots+j_{N}=i}\prod_{l=1}^{N}{x_{l}^{j_{l}}}.$
We denote $\omega_{m}=\prod_{1\leq i<j\leq N}(x_{i}^{(m)}-x_{j}^{(m)})$ where $x^{(m)}_{i}$ are Chern roots of line bundles defined in (40).
Consider the equivariant cohomology ring of $Fl$. Let $\omega=\prod_{k=1}^{n}\omega_{k}=\prod_{k=1}^{n}\prod_{1\leq i<j\leq N_{k}}(x^{(k)}_{i}-x^{(k)}_{j})$. By the short exact sequence (37), the equivariant cohomology ring of $Fl$ is
$H^{*}_{S}(Fl)\cong\mathbb{Q}[\{x^{(i)}_{j}\}_{i=1,\cdots,n+1,j=1,\cdots,N_{i}}]^{W}/I$
where
$I=\{f\in\mathbb{Q}[\lambda_{1},\ldots,\lambda_{N_{n+1}}][\{x^{(i)}_{j}\}_{i=1,\cdots,n+1,j=1,\cdots,N_{i}}]^{W}\big{|}\,f\cup\omega\in I^{ab}\}.$
@@
(46) $\widetilde{c_{t}^{S}(S_{k})}=\prod_{j=1}^{N_{k}}(t-x^{(k)}_{j})=\sum_{l=0}^{N_{k}}t^{N_{k}-l}e_{l}(-\vec{x}^{(k)}).$
In $H^{*}_{S}(Fl)$, the Chern polynomial of the quotient bundle $S_{l}/S_{k}$ which is defined in (34) satisfies the following relation
(47) $c_{t}^{S}(S_{k})\cup c_{t}^{S}(S_{l}/S_{k})=c_{t}^{S}(S_{l}).$
@@
(48) $\prod_{j=1}^{N_{l}}\left(x^{(k)}_{i}-x^{(l)}_{j}\right)\prod_{m=k+1}^{l-1}\omega_{m}=0,\,\,i=1,2,\ldots,N_{k}.$
@@
We use an induction argument on the difference $l-k$. For $l-k=1$, the equation (48) holds by (41). We assume (48) holds for $l-k\leq p-1$. For $l-k=p$, we consider
(49) $A:=\det\begin{bmatrix}\prod_{j=1}^{N_{l}}\left(x^{(k)}_{i}-x^{(l)}_{j}\right)&\prod_{j=1}^{N_{l}}\left(x_{1}^{(l-1)}-x^{(l)}_{j}\right)&\cdots&\prod_{j=1}^{N_{l}}\left(x_{N_{l-1}}^{(l-1)}-x^{(l)}_{j}\right)\\ \left(x_{i}^{(k)}\right)^{N_{l-1}-1}&\left(x_{1}^{(l-1)}\right)^{N_{l-1}-1}&\cdots&\left(x_{N_{l-1}}^{(l-1)}\right)^{N_{l-1}-1}\\ \vdots&\vdots&\ddots&\vdots\\ x_{i}^{(k)}&x_{1}^{(l-1)}&\cdots&x_{N_{l-1}}^{(l-1)}\\ 1&1&\cdots&1\\ \end{bmatrix}\cdot\prod_{m=k+1}^{l-2}\omega_{m}$
@@
$A=\sum_{a=1}^{N_{l}}e_{a}(-\vec{x}^{(l)})h_{N_{l}-N_{l-1}-a}(x^{(k)}_{i},\vec{x}^{(l-1)})\prod_{j=1}^{N_{l-1}}\left(x^{(k)}_{i}-x^{(l-1)}_{j}\right)\prod_{m=k+1}^{l-1}\omega_{m}$
@@
On the other hand, expanding (49) along the first row, by (41), (44) we have
$A=\prod_{j=1}^{N_{l}}\left(x^{(k)}_{i}-x^{(l)}_{j}\right)\prod_{m=k+1}^{l-1}\omega_{m}$
So (48) holds for $l-k=p$. ∎
@@
(50) $\widetilde{c_{t}^{S}(S_{l}/S_{k})}=\sum_{a=0}^{N_{l}-N_{k}}t^{a}\left[\sum_{b=0}^{N_{l}-N_{k}-a}e_{b}(-\vec{x}^{(l)})h_{N_{l}-N_{k}-a-b}(\vec{x}^{(k)})\right].$
@@
By (37) and (47), it suffices to check
$\widetilde{c_{t}^{S}(S_{k})}\cup\sum_{a=0}^{N_{l}-N_{k}}t^{a}\left[\sum_{b=0}^{N_{l}-N_{k}-a}e_{b}(-\vec{x}^{(l)})h_{N_{l}-N_{k}-a-b}(\vec{x}^{(k)})\right]\cup\omega=\widetilde{c_{t}^{S}(S_{l})}\cup\omega.$